OpenStudy (samigupta8):

if |(x^2+2x+2)+(3x+7)| < |x^2+2x+2| + |3x+7|. then set of complete values of x is

OpenStudy (anonymous):

heck of a question !

OpenStudy (anonymous):

You can use desmos.com/calculator and do it yourself ;)

OpenStudy (anonymous):

I'll go the human way, :P , ok give me a moment.

OpenStudy (abb0t):

Anyone can use a calculator @koalamon but that doesn't mean the person understands what they're doing. It also doesn't teach them the fundamentals for more advance mathematics courses. It's best if the person knows what they're doing. Good job @kappa007 for tackling the problem head on.

OpenStudy (anonymous):

k, just a sec, someone's in urgent need. I'll be back in like 5 mins. I hope you can wait.

OpenStudy (samigupta8):

okk...

OpenStudy (anonymous):

He didn't ask for anyone to explain, he asked to find the complete values of x, and the calculator requires you to put in the information in an understandable format. If it comes out as a graph you will need to isolate anywhere where y=0

OpenStudy (anonymous):

okay, the quadratic expression here has a negative discriminant, do you agree ?

OpenStudy (anonymous):

no ?

OpenStudy (anonymous):

So, if the discriminant is negative, it means the quadratic expression doesn't cut the x-axis at all i.e, it is always positive or always negative.

OpenStudy (samigupta8):

ryt...

OpenStudy (anonymous):

We can take 2 cases when \[x \ge \frac{ -3 }{ 7 }\] and \[x < \frac{ -3 }{ 7 }\]

OpenStudy (anonymous):

Try it out. See what happens. Tell me if there's doubt.

OpenStudy (ankit042):

I am not completely sure but how about using this |A+B| < |A| + |B| so either |A| <0 or |B| < 0 or both |A| and |B| <0 now if solve these inequality we reach our answer.

OpenStudy (anonymous):

In this case, A remains always positive, so that's a help.

OpenStudy (samigupta8):

kappa007 how are u getting it as -3/7 i m getting it -7/3

OpenStudy (ankit042):

yeah true we can write A as (x+1)^2+1...and @samigupta8 I also feel it should be -7/3

OpenStudy (anonymous):

ohh , silly mistake, ignore it, you're right

OpenStudy (anonymous):

that's obv -7/3

OpenStudy (anonymous):

did you get the ans ?

OpenStudy (samigupta8):

okk...so now we will solve it like for x<-7/3 |(x^2+2x+2)+(3x+7)| < x^2+2x+2-3x-7 (x^2+2x+2)+(3x+7) < x^2-x-5 x^2+5x+9<x^-x-5 6x<-14 x<-7/3

OpenStudy (samigupta8):

m i rytt...

OpenStudy (anonymous):

so, that's you answer. When we assume the 2nd case, the result is in agreement. That itself is the required interval.

OpenStudy (anonymous):

yup

OpenStudy (samigupta8):

and for case 1st it will all come out to be zero

OpenStudy (anonymous):

one point, you can directly cancel out the quadratic exp on both sides before simplifying.

OpenStudy (anonymous):

for case 1, it comes out 0 < 0 which is absurd, so we infer that the interval doesn't have any solution.

OpenStudy (samigupta8):

okk..thankuu...